# ZERO RESOURCE CODE-SWITCHED SPEECH BENCHMARK USING SPEECH UTTERANCE PAIRS FOR MULTIPLE SPOKEN LANGUAGES

### 17 Dec 2023 v2

@@ 3.1 Data generation and validation
As suggested in [14], we conducted human validation by multiple bilingual speakers. Each human annotator was required to label whether the paired sentences were valid, meaning that the presumed correct sentence in each pair should (1) actually make sense and be meaningful and grammatically acceptable, (2) be indeed better than the presumed wrong one in the aspects above. Pairs failing to meet the above two requirements would be labeled invalid. To ensure the annotation quality, the hired annotators were required to complete an annotation trial on some sampled paired sentence data with pre-defined ground truths. Human annotators were required to get at least $95\%$ accuracy before proceeding to the data annotating process. A pair of correct and wrong sentences was included in the task if most annotators labeled it valid.
@@ 3.2 Code-switched data statistics
This zero resource code-switched speech benchmark has three tracks based on three code-switched language pairs, including Spanish-English (es-en), French-English (fr-en), and Chinese-English (zh-en), with 7263, 4020, and 3176 human-validated data samples, respectively. For each language pair, all available bilingual speaker configurations were adopted from the Amazon Polly text-to-speech system to synthesize each utterance. All the synthesized utterances had a sample rate of 22.5kHz originally and were later resampled to 16kHz to match the configurations of the speech encoders.
@@ 5 Results
Please note that although we tried to restrict the length difference of the sentences to balance the length of the synthesized utterances between the correct and wrong versions, it is still impossible to make the lengths exactly matched for every pair. Thus, the use of direct likelihood comparisons may lead to a bias in favor of the shorter sentence, which was generally the wrong one. While most of the speech-based baselines were not significantly influenced, perhaps because of their informative units, and could apparently distinguish the two utterances as the span-PP scores of the two utterances differed a lot, the random baseline was heavily misled since its units were randomly assigned, making its performance fall below 50%.
@@ 5.1 Multilingual pre-training
As for the zh-en track, except for XLS-R 1B, all the models that included Chinese in their pre-training slightly outperform their monolingual counterpart (Wav2Vec2.0 Large), though the differences are insignificant. This may be a result of relatively inadequate pre-training data in Chinese compared with the Spanish and French pre-training data. For mHuBERT, the performance on the zh-en track is quite close to that of HuBERT Base since Chinese speech data were absent during its pre-training stage.
@@ 5.2 Model size and pre-training languages
Next, we find that the system with XLS-R 0.3B significantly outperforms that with XLSR-53, which may imply that the multilingual pre-training with broader coverage of languages provides better generalizability for code-switched speech and thus induces better performance on our benchmark.

### 18 Mar 2024 v3

@@ 3.1 Data generation and validation
As suggested in [14], we conducted human validation by multiple bilingual speakers. Each human annotator was required to label whether the paired sentences were valid, meaning that the presumed correct sentence in each pair should (1) actually make sense and be meaningful and grammatically acceptable, (2) be indeed better than the presumed wrong one in the aspects above. Pairs failing to meet the above two requirements would be labeled invalid. To ensure the annotation quality and consistency, the hired annotators were required to complete an annotation trial on some sampled paired sentence data with pre-defined ground truths. Human annotators were required to get at least $95\%$ accuracy before proceeding to the data annotating process. A pair of correct and wrong sentences was included in the task if most of the annotators labeled it valid.
@@ 3.2 Code-switched data statistics
This zero resource code-switched speech benchmark has three tracks based on three code-switched language pairs, including Spanish-English (es-en), French-English (fr-en), and Chinese-English (zh-en), with 7263, 4020, and 3176 human-validated data samples, respectively. For each language pair, all available bilingual speaker configurations were adopted from the Amazon Polly text-to-speech system to synthesize utterances, and the configurations of the two utterances in each pair were the same. All the synthesized utterances had a sample rate of 22.5kHz originally and were later resampled to 16kHz to match the configurations of the speech encoders.
@@ 5 Results
Please note that although we tried our best to restrict the length difference of the sentences to balance the length of the utterances between the correct and wrong versions, it’s still impossible to make the lengths exactly matched for every pair. Thus, there may be a bias in favor of the shorter utterance, which was generally the wrong one. While most of the speech-based baselines were not significantly influenced, perhaps because of their informative units, and could apparently distinguish the two utterances as the span-PP scores of the two utterances differed a lot, the random baseline was heavily misled since its units were randomly assigned, making its performance fall below 50%. Nevertheless, this only shifts the reference level for comparison, and the interpretation of the results remains the same.
@@ 5.1 Multilingual pre-training
As for the zh-en track, except for XLS-R 1B, all the models that included Chinese during pre-training slightly outperform their monolingual counterpart (Wav2Vec2.0 Large), though the differences are insignificant. This may be due to relatively inadequate pre-training data in Chinese compared with those in Spanish and French. For mHuBERT, the performance on the zh-en track is close to that of HuBERT Base since Chinese was absent during pre-training.
@@ 5.2 Model size and pre-training languages
Next, we find that the system with XLS-R 0.3B significantly outperforms that with XLSR-53, which may imply that the multilingual pre-training with broader coverage of languages provides better generalizability for code-switching and thus induces better performance on our benchmark.
@@ 6 Conclusion
7 Acknowledgement
We thank the National Center for High-performance Computing (NCHC) of the National Applied Research Laboratories (NARLabs) in Taiwan for providing computational and storage resources.